solve this... k over 9 = 10 over 18
Found 3 solutions by rfer, rosey51235, solver91311:
Answer by rfer(16322)

k/9=10/18
18k=90
k=90/18
k=5

Answer by rosey51235(30)

There's a whole big complicated proportion for this, or I could just tell you that since 9 and 18 are both on the bottom, since 9 is half of 18, whatever is over 18 will be in half over 9.
so if 10 is over 18 then 5 should be over 9.
But if you wanted to do proportions, you would cross multiply, end up with 18k=90, divide by 18 on both sides and still get 5.
Hope I helped! :)
-Rosey

Answer by solver91311(24713)



Cross-multiply and then solve the resulting equation in k.


John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism